# Steering LLM Reasoning Through Bias-Only Adaptation

### v1

@@ Appendix D LoRA
To let the offset be token-specific, we replace the fixed steering vector with a low-rank adaptor (LoRA) (Hu et al., 2022) applied to the MLP down-projection in every transformer layer:
@@
Appendix E Logit Lens. GPT Prompt
GPT Prompt for Token Clustering {minted} [linenos]text You will be given a list of tokens together with a score. You should translate all non-english tokens and suggest the main topics that unite the biggest subsets of tokens in the list. ¡list¿
Appendix F Computational Resources
All models were trained on $16$ H100 GPUs. Qwen2.5-1.5B models were trained for approx. $9$ hours, Qwen2.5-Math-1.5B for approx. $2.5$ hours, llama3.1-8b-instruct for approx. $9$ hours, llama3.1-8b-instruct for approx. $120$ hours.

### v2

@@ Appendix D LoRA
To let the offset be token-specific, we replace the fixed steering vector with a low-rank adaptor (LoRA) Hu et al. (2022) applied to the MLP down-projection in every transformer layer:
@@
Appendix E Steering Vectors on GSM8K and MATH
Experimental Setup.
We conduct experiments on four pretrained transformer checkpoints: Qwen-2.5-1.5B (Team, 2024), Qwen-2.5-Math-1.5B (Yang et al., 2024), Llama-3.1-8B, and Llama-3.1-8B-Instruct Grattafiori et al. (2024). For each model, we evaluate three training regimes: (i) full fine-tuning, (ii) training only steering vectors, and (iii) training only LoRA (Hu et al., 2022) adapters which may be viewed as adaptive steering vectors (Appendix D, Mack and Turner (2024)). For LoRAs we use rank $4$. In the latter two cases, all other parameters are kept frozen.
We use two mathematical datasets for training and evaluation. The GSM8K training split contains 8,790 problems Cobbe et al. (2021); for evaluation, we randomly subsample 500 items from its original split to shorten iteration time. The MATH corpus Hendrycks et al. (2021b) provides 12,000 training examples, and its evaluation split consists of 500 items, which we use as is. We report $\text{mean}@8=\mathbb{E}_{x}[\mathbb{E}_{i=0}^{8}\ r(x,y)]$ win rates on each dataset.
All experiments are implemented using the transformers library (Wolf et al., 2019) and the vllm inference engine Kwon et al. (2023). Additional hyperparameter details are provided in Appendix I.
Result.
Table˜4 summarizes accuracy for the pretrained model, fully-tuned models, steering vectors, and LoRA adapters. As expected, RL training yields large gains on mathematical benchmarks on all setups. Steering vectors achieve similar improvements across nearly all model-dataset pairs and even exceed full fine-tuning in some cases (e.g., Qwen2.5-1.5B when evaluated on MATH-500 and LLaMa3.1-8B when trained on GSM8K and evaluated on MATH-500, both being base models that did not undergo instruction tuning), which we attribute to the implicit regularization of updating far fewer parameters.
If we accept the working assumption that a steering vector can only amplify features that the original network already contains and cannot create new ones, the table gives direct evidence for that view: when the base model "knows" how to solve the task, steering is usually enough to reach the same quality as full fine-tuning.
Appendix F Logit Lens. GPT Prompt
GPT Prompt for Token Clustering You will be given a list of tokens together with a score. You should translate all non-english tokens and suggest the main topics that unite the biggest subsets of tokens in the $\par<list$>
Appendix G Training Cost Savings
(a) Number of Parameters
Qwen2.5-7B Llama3.1-8B
Full-Tune 7.6 B 8 B
Steering 100 K 131 K
(b) Optimizer Memory
Qwen2.5-7B Llama3.1-8B
Full-Tune 7.1 GB 7.5 GB
Steering 98 KB 128 KB
(c) Per-step Time
Qwen2.5-7B Llama3.1-8B
Full-Tune 5.30 s 5.32 s
Steering 0.06 s 0.07 s
(d) Overall Time (314 steps $\approx$ 1 epoch)
Qwen2.5-7B Llama3.1-8B
Full-Tune 27.7 m 27.8 m
Steering 0.314 m 0.36 m
Table 5: Resource-efficiency comparison of full fine-tuning versus steering across three model sizes.
Appendix H Computational Resources
All models were trained on $16$ H100 GPUs. Qwen2.5-1.5B models were trained for approx. $9$ hours, Qwen2.5-Math-1.5B for approx. $2.5$ hours, LLaMa3.1-8B-It for approx. $9$ hours, LLaMa3.1-8B-It for approx. $120$ hours.
Appendix I Hyperparameters
$\mathrm{lr}$ $\mathrm{num\_generations}$ Qwen-2.5-1.5B GSM8K Full-Tune $2\times 10^{-5}$ 64 Steering $5\times 10^{-4}$ 64 LoRA-1 $5\times 10^{-4}$ 64 LoRA-4 $5\times 10^{-4}$ 64 MATH Full-Tune $2\times 10^{-5}$ 64 Steering $5\times 10^{-4}$ 64 LoRA-1 $5\times 10^{-4}$ 64 LoRA-4 $5\times 10^{-4}$ 64 Qwen-2.5-Math-1.5B GSM8K Full-Tune $2\times 10^{-5}$ 16 Steering $1\times 10^{-3}$ 16 LoRA-1 $5\times 10^{-4}$ 16 LoRA-4 $5\times 10^{-4}$ 16 MATH Full-Tune $2\times 10^{-5}$ 16 Steering $1\times 10^{-3}$ 16 LoRA-1 $5\times 10^{-4}$ 16 LoRA-4 $5\times 10^{-4}$ 16 Llama-3.1-8B GSM8K Full-Tune $5\times 10^{-6}$ 64 Steering $5\times 10^{-4}$ 64 LoRA-1 $1\times 10^{-4}$ 64 LoRA-4 $1\times 10^{-4}$ 64 MATH Full-Tune $5\times 10^{-6}$ 64 Steering $5\times 10^{-4}$ 64 LoRA-1 $1\times 10^{-4}$ 64 LoRA-4 $1\times 10^{-4}$ 64 Llama-3.1-8B-Instruct GSM8K Full-Tune $1\times 10^{-6}$ 16 Steering $2\times 10^{-4}$ 16 LoRA-1 $6\times 10^{-4}$ 16 LoRA-4 $1\times 10^{-4}$ 16 MATH Full-Tune $1\times 10^{-6}$ 16 Steering $2\times 10^{-4}$ 16 LoRA-1 $3\times 10^{-4}$ 16 LoRA-4 $3\times 10^{-4}$ 16
Table 6: Hyperparameter settings for each model and training setup.